The greatest acceleration or deacceleration train may have is d find  minimum time in which train may reach from one station to another station  separated by a distance d is 

Grade:11

1 Answers

Vikas TU
7 years ago
d = ut + 0.5dt^2
Differentiate t w.r.t d then,
1 = udt/dd +d*t*dt/dd
dt/dd(u+t*d) = 1
dt/dd = 1/(u + t*d)
1/(u+td) = 0 
when t tends to infinity.